I have a graph with numbers 10,30,60,100,150,210,280,360 and so on. I am having trouble finding the correct fomular for this problem.
thank you for your help

Answer by venugopalramana(3286)

I have a graph with numbers 10,30,60,100,150,210,280,360 and so on. I am having trouble finding the correct fomular for this problem.
thank you for your help
WHAT DO YOU WANT?
FORMULA FOR N TH. TERM???ASSUMING SO ......NOTE THE FOLLOWING
A2-A1=30-10=20
A3-A2=60-30=30
A4-A3=100-60=40
A5-A4=150-100=50
A6-A5=210-150=60
A7-A6=280-210=70
A8-A7=360-280=80
..............
WE ARE HAVING A.P. ON RHS 20,30,40,50...ETC..WITH D=10...TN=20+(N-1)*10=10N+10=10(N+1)...HENCE...TN-1=10(N-1+1)=10N
AN-A(N-1)=10N.........CHECK ABOVE PUTTING N=2,3,4 ETC....
------------------------ADDING ALL
AN-A1=20+30+................10N
AN-10=(N-1)[20+10N]/2=5(N-1)(N+2)
AN=10+5(N-1)(N+2)
CHECK
N=1.......A1=10
N=2.......A2=10+5*4=30
N=3.......A3=10+5*2*5=60
N=4.......A4=10+5*3*6=100.....
......ETC....OK